University of Vermont

UVM Blogs

fcs

  • 22 May 2015: Friday

    fcs on Frank's Activity Log
    LDAP: No issues in the nightly LDAP update. No issues in the morning ACCOUNT update. Backups: The usual three failures for this week. bujbod2: The trial run was a success.  3.2TB written there […]
    Posted 12 hours ago.
  • 21 May 2015: Thursday

    fcs on Frank's Activity Log
    LDAP: Nightly LDAP update issues: one – a new student employee, found and merged. No daily ACCOUNT update issues. Backups: CoM/IS and disk clones continue. Special processing for old AIX box […]
    Posted yesterday.
  • 20 May 2015: Wednesday

    fcs on Frank's Activity Log
    LDAP: Nightly LDAP update: 1 issue – a student, employee – found, merged. Daily ACCOUNT update: no issues Backups: CoM/IS and Disk clones continue. SB backups still having a failure.  Pinging the […]
    Posted 2 days ago.
  • 19 May 2015: Tuesday

    fcs on Frank's Activity Log
    LDAP: Two issues in the LDAP update:  Two new student employees.   Found their student record and merged the employee info with it. No issues in the morning ACCOUNT update. Backups: That CoM/IS system […]
    Posted 3 days ago.
  • 18 May 2015: Monday

    fcs on Frank's Activity Log
    LDAP: One issue over the weekend: Found the student and merged the new employee info in. No problems in the daily ACCOUNT updates. Backups: Unexpected backup failures – admins notified. bujbod2 […]
    Posted 4 days ago.
  • 15 May 2015: Friday

    fcs on Frank's Activity Log
    LDAP: Patched up three issues from Wednesday night and modifications to those same entries from Thursday night in the LDAP update issues field. No issues in the ACCOUNT update the past two mornings. Update […]
    Posted 1 weeks ago.
  • 13 May 2015: Wednesday

    fcs on Frank's Activity Log
    LDAP: An “Updating the WRONG entry” issue in the nightly LDAP update: Banner adds an SSN to an entry and we find it matches an employee. Student account not activated. Entries merged; Account Services […]
    Posted 2 weeks ago.
  • 12 May 2015: Tuesday

    fcs on Frank's Activity Log
    LDAP: Two issues in LDAP update – found and merged. No issues in ACCOUNT update. Backups: CoM/IS full from the weekend continues. Clones are being delayed so I can install the nsrexecd hotfix when […]
    Posted 2 weeks ago.
  • 11 May 2015: Monday

    fcs on Frank's Activity Log
    LDAP: Nine ldap update issues to deal with from while I was out on vacation. April 28: A single item match (that was partially handled) – completed. A Graduate Student becomes an employee.  found and […]
    Posted 2 weeks ago.
  • 24 April 2015: Friday

    fcs on Frank's Activity Log
    LDAP: No issues in last night’s LDAP update. No issues in this morning’s ACCOUNT update. Backups: Just the one error that we are ignoring because we are replacing that system. Investigating why […]
    Posted 4 weeks ago.
  • 23 April 2015: Thursday

    fcs on Frank's Activity Log
    LDAP: No issues in last night’s LDAP update. No issues in this morning’s ACCOUNT update. Backups: No unexpected failures overnight. NW seems to be losing even more ssids on disk which it knows […]
    Posted 5 weeks ago.
  • 22 April 2015: Wednesday

    fcs on Frank's Activity Log
    LDAP: No issues in last night’s LDAP update. No issues in this morning’s ACCOUNT update. Backups: Clones finished up last night. windowbox set up RHEL7: fighting with selinux and nrpe on bujbod2
    Posted 5 weeks ago.
  • 21 April 2015: Tuesday

    fcs on Frank's Activity Log
    LDAP: The Giraffe migration broke the delivery of feed files to giraffe last night.  Spent two hours (9pm-11pm) working with Kent and Mike to get it fixed. Mike updated the LDAP ACLs because of the giraffe […]
    Posted last month.
  • 20 April 2015: Monday

    fcs on Frank's Activity Log
    LDAP: No issues in the nightly LDAP update over the weekend. No issues in the morning ACCOUNT update over the weekend. Backups: DD SR 70755774 – someone finally called back at 4:15pm (after I had left […]
    Posted last month.
  • 17 April 2015: Friday

    fcs on Frank's Activity Log
    LDAP: No issues in the LDAP update last night. No issues in the ACCOUNT update this morning. Backups: Thirty (30) windows systems refuse to let our networker server back them up this morning.  This […]
    Posted last month.
  • 16 April 2015: Thursday

    fcs on Frank's Activity Log
    LDAP: No issues in LDAP update. No issues in ACCOUNT update. Backups: Disk clones finished up at 3am this morning. A change I made to start removing DD690 attachments yesterday caused the bootstrap […]
    Posted last month.
  • 15 April 2015: Wednesday

    fcs on Frank's Activity Log
    LDAP: Automated Purge Day: four errors (email bounces that already exist) No issues in the nightly LDAP update. No issues in the morning ACCOUNT update. Backups: Disk clones continue, with just 5TB […]
    Posted last month.
  • 14 April 2015: Tuesday

    fcs on Frank's Activity Log
    LDAP: Monday night’s update had a serious problem: OSP sent a birthday that started with a zero (0), which happens to be illegal in the LDAP schema for integers.  The issue was paged at 8:30pm.  I was paged at […]
    Posted last month.
  • 13 April 2015: Monday

    fcs on Frank's Activity Log
    LDAP: No issues in the nightly LDAP update over the weekend. No issues in the daily ACCOUNT update over the weekend. Backups: Tape Movement: 190 tapes went to the vault, 240 came back. Reclaimed […]
    Posted last month.
  • 10 April 2015: Friday

    fcs on Frank's Activity Log
    LDAP: No issues in the nightly update the past two nights. No issues in the daily Account update the past two mornings. Backups: Weekly clone process completed. More backup clients […]
    Posted last month.
Skip to toolbar